OpenStudy (anonymous):

Find the distance between the given points. A(5, 3) and B(7, -4) Distance = √ ?

OpenStudy (anonymous):

sqrt( (x-x0)^2+(y-y0)^2)

OpenStudy (anonymous):

think about it in terms of triangles

OpenStudy (anonymous):

Distance = \[\sqrt{(7-5)^2 +(-4-3)^2 } ..... = \sqrt{(2)^2 + (-7)^2} ... = \sqrt{4+49} = \sqrt{53} \]